OpenStudy (anonymous):

integrate sqrt((e^t) − 3) using u substitution

OpenStudy (luigi0210):

Hello @ginsuman Welcome to Openstudy :)

OpenStudy (luigi0210):

\[\int\limits \sqrt{e^t-3} dx\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Do you know how to start this off?

OpenStudy (anonymous):

well....i believe i want to find u and u' first....

OpenStudy (anonymous):

so u = e^t-3 and u' = 1/2(e^t-3)^-1/2(e^t) (not 100% of the derivative tho)

OpenStudy (luigi0210):

The derivative of e^x is e^x

OpenStudy (anonymous):

right. but...what is the derivative of sqrt(e^t-3) dt ?

OpenStudy (luigi0210):

If you ask me it'd be easier to set u=e^x-3

OpenStudy (luigi0210):

Wouldn't you agree? Then we'd just be left with: \[\int\limits \sqrt{u} dx\]

OpenStudy (anonymous):

yes...but then what inverse trig function is used to complete the integral?

OpenStudy (luigi0210):

u=e^x-3 du=e^x dx

OpenStudy (anonymous):

i was wrong earlier....u must be equal to the original problem - not what was inside the sqrt

OpenStudy (luigi0210):

Oh, then in that case you were right. \[u= \sqrt{e^x-3}\] \[du=\frac{1}{2}(e^x-3)^{-1/2}*e^x\]

OpenStudy (anonymous):

yes....that's what i had originally. now....using inverse trig functions....how do I proceed? i'm not sure how to recognize which inverse trig function to use.

OpenStudy (luigi0210):

I don't think there are any inverse trig values involved. But then again I haven't taken calculus in over a year.

OpenStudy (anonymous):

there is definitely an inverse trig function involved. i have the answer (from a textbook) but I am trying to figure out the detailed step-by-step solution on how to get the answer the book gives.

OpenStudy (luigi0210):

Well let's see the answer then.

OpenStudy (luigi0210):

Seeing the answer will help.

OpenStudy (anonymous):

ok....hold on...

OpenStudy (anonymous):

i'm going to try using the Equation editor...

OpenStudy (anonymous):

\[2(\sqrt(e^t-3))-2(\sqrt(3))\arctan(\sqrt(e^t-3)/\sqrt(3))+C\]

OpenStudy (luigi0210):

Okay I see, hold on

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

are you still there?

OpenStudy (luigi0210):

Yea, do you still need help?

OpenStudy (anonymous):

yes please. although i have the answer i do not understand how to get there. i'm interesting in learning the step-by-step process

OpenStudy (luigi0210):

Oops, sorry it's in respect to t.

OpenStudy (luigi0210):

He provided the answer too.

OpenStudy (luigi0210):

I honestly have no idea, I've been calling out to all usersand so far no luck.

OpenStudy (anonymous):

\[Let I=\int\limits \sqrt{e ^{x}-3}dx\] \[put \sqrt{e ^{x}-3}=t,e ^{x}-3=t ^{2},e ^{x}=t ^{2}+3,e ^{x}dx=2tdt\] \[\left( t ^{2}+3 \right)dx=2t dt ,dx=\frac{ 2tdt }{t ^{2}+3 }\]

OpenStudy (anonymous):

\[I=\int\limits \frac{ 2t ^{2}dt }{t ^{2}+3 }=2 I1+c\]

OpenStudy (anonymous):

\[I1=\int\limits \frac{ t ^{2}+3 }{t ^{2}+3 }dt-\int\limits \frac{ 3dt }{ t ^{2}+3 }\] \[I1=t-3 \frac{ 1 }{ \sqrt{3} }\tan^{-1} \frac{ t }{ \sqrt{3} }\]

OpenStudy (anonymous):

I think now you can solve.